# A Novel Garment Transfer Method Supervised by Distilled Knowledge of Virtual Try-on Model

### v1

@@ 4.3.2 Garment Transfer Method
Baseline We compare our method with other state-of-the-art garment transfer methods, which consist of SwapNet (Raj et al., 2018), CT-Net (Yang and Lin, 2021), and PASTA-GAN (Xie et al., 2021). We trained SwapNet on the Zalando dataset with an unofficial implementation [^4] https://github.com/andrewjong/SwapNet. CT-Net requires paired images of the same person wearing the same garment in different poses for training. Therefore, we utilized its official implementation and pre-trained checkpoint on Deepfashion2 (Ge et al., 2019) and test it on the Zalando dataset. In addition, PASTA-GAN adopts its official implementation and pre-trained checkpoint.
@@
Table 6: Quantitative comparison with garment transfer methods in garment transfer from ${{{\cal I}_{B}}}$ to ${{{\cal I}_{A}}}$.
Ours SwapNet CT-Net PASTA-GAN
IS↑ 2.75 2.73 2.48 2.71
hyperIQA↑ 35.94 27.20 33.24 32.93
Table 7: Quantitative comparison with garment transfer methods in garment transfer from ~${{\tilde{\cal I}_{AB}}}$ to ${{{\cal I}_{B}}}$.
MSE↓ PSNR↑ SSIM↑ FID↓ LPIPS↓ IS↑
Ours 389.029 23.57 0.863 19.14 0.108 2.734
SwapNet 1705.66 16.31 0.677 67.45 0.269 2.557
CT-Net 2980.83 13.86 0.626 86.69 0.352 2.266
PASTA-GAN 621.76 21.07 0.845 15.03 0.105 2.725
SwapNet learns category embedding to infer pixel content, which places a significant learning burden on the CNN model. When the garment being transferred has intricate texture and local details, SwapNet tends to blur textures and discard high-frequency details, resulting in unrealistic outcomes, particularly in face and garment regions. This leads to a decrease in hyperIQA by 24.31%. CT-Net’s robustness on the Zalando dataset is low without re-training it, and its training strategy falls under self-supervision learning, which causes differences in input between the training and testing phases. Moreover, the arm inferring mechanism of CT-Net is implicit, making it challenging to handle the cases of exposing new skin as the first and second rows in Fig. 22. PASTA-GAN disentangles the garment with the pose by dividing the garment image into several patches, but the semantics between patches are discontinuous when reconstructing the garment structure as the target pose. This creates an apparent boundary between patches, leading to unrealistic transfer results, with SSIM decreasing by 10.61%. In contrast, our method warps the garment instead of mapping it, has an explicit arm regrowth mechanism, and attains the highest scores of IS and hyperIQA among the garment transfer methods discussed.
As Fig. 23 shows, when the transferred garment is in an extreme pose, the garment warping continuity makes it challenging to tear down the cover region and leave a blank, and the sleeve of the transferred garment is put in an incorrect position. To avoid transfer failure, the user should not leverage such a person image as garment condition. As unobserved content inference is still a challenging topic in the community, our future work will involve combining warping with rendering to handle the transferred garment occlusion.
Figure 23: Failure cases.
@@ 5 Conclusion
This paper proposes a novel garment transfer method supervised with knowledge distillation from virtual try-on. Our method first reasons the transfer parsing to provide shape prior for downstream tasks. We employ a multi-phase teaching strategy to supervise the transfer parsing reasoning model during training. When transferring a garment to another person, it learns the response and feature knowledge from the try-on parsing reasoning model. When transferring the garment back to its owner, it absorbs the hard knowledge from the ground truth. Guided by the transfer parsing, we map the position of the transferred garment mask toward the target shape to prevent distortion. Afterward, we estimate a progressive flow for precise warping of the garment by learning shape and content correspondence. To ensure that the boundary and internal structure of the garment are changed reasonably, we supervise the training of the flow warping model using target shape and warping knowledge from virtual try-on. To better preserve body features in the transfer result, we propose a well-designed training strategy for the arm regrowth task, which propagates the source arm part and reasons the new arm part. Experiments demonstrate that our method has state-of-the-art performance compared with other virtual try-on and garment transfer methods in garment transfer, especially for preserving garment texture and body features.

### v2

@@ 4.3.2 Garment Transfer Method
Baseline We compare our method with other state-of-the-art garment transfer methods, which consist of SwapNet [3], CT-Net [4], and PASTA-GAN [13]. CT-Net requires paired images of the same person wearing the same garment in different poses for training. Therefore, we utilized its official implementation and pre-trained checkpoint on Deepfashion2 [37] and test it on the Zalando dataset.
@@
[…]
Method ${{{\cal I}_{B}}\to{{\cal I}_{A}}}$${{\tilde{\cal I}_{AB}}\to{{\cal I}_{B}}}$
IS↑ hyperIQA↑ MSE↓ PSNR↑ SSIM↑ FID↓ LPIPS↓ IS↑
Ours 2.75 35.94 389.029 23.57 0.863 19.14 0.108 2.734
[…]
CT-Net 2.48 33.24 2980.83 13.86 0.626 86.69 0.352 2.266
PASTA-GAN 2.71 32.93 621.76 21.07 0.845 15.03 0.105 2.725
SwapNet learns statistical embedding for each category and paints the target shape at a feature level, imposing a heavy learning burden on the CNN model. As shown in the first row of Fig. 22, when handling in local region with complex textures or structures, like faces, SwapNet is prone to miss high-frequency details and blur texture, where hyperIQA is decreased by 24.31%. CT-Net treats the garment transfer as the pose transfer task and implements it in a self-supervision manner. As shown in the second row of Fig. 22, in case of no retraining on the Zalando dataset, CT-Net faces the generation and adaption problems in the testing phase, where the pose feature is misaligned and arm is not exposed in transferring a garment with a different type. PASTA-GAN disentangles pose features from garments via patch division and trains a garment transfer model to re-entangle them in a self-supervision manner. As shown in the third and fourth rows in Fig. 22, patch division leaves apparent boundaries and makes the re-entangled garment discontinuous across patches, decreasing SSIM by 10.61%. In contrast, our method leverages a warping mechanism to change the pose and handle arm exposure via an explicit regrowth mechanism, achieving the highest scores of IS and hyperIQA among the garment transfer methods discussed.
4.4 Failure Case
Our method has limited performance in transferring a garment with the self-occlusion phenomenon. As Fig. 23 shows, this phenomenon easily appears when a person crosses or raises her arms. With the continuity property in our garment warping, it is hard to tear down the self-occlusion region and leave a blank, thus, the sleeve is prone to be put in an incorrect position. To avoid these failure cases, a straightforward trick is to not take these self-occlusion images as inputs, and for future technical progress, we will focus on establishing the explicit mapping relationship of garments with different poses, via the 3D dense pose or human body model, and explore to employ inpainting mechanism to infer the unobserved region.
Figure 23: Failure cases of our method.
4.5 Generalization
To verify the generalization capacity, we evaluate our method on VITON-HD [38] and deepfashion2 [37] datasets in case of not retraining our method on them. The qualitative examples and quantitative results are shown in Fig. 24 and Table 6. Experimental results show that our method still can robustly generate high-quality transferred images. It demonstrates that transfer parsing reasoning and garment warping mechanisms are available for general cases, and our method can be applied in real transfer scenes.
Figure 24: Qualitative examples of our method on VITON-HD and Deepfashion2 datasets.
Table 6: Quantitative results of our method VITON-HD and Deepfashion2 datasets.
IS↑ hyperIQA↑
VITON-HD 2.98 38.32
Deepfashion2 2.56 40.06
@@ 5 Conclusion
This paper proposes a novel garment transfer method supervised with knowledge distillation from virtual try-on. Our method first reasons the transfer parsing to provide shape prior to downstream tasks. We employ a multi-phase teaching strategy to supervise the training of the transfer parsing reasoning model, learning the response and feature knowledge from the try-on parsing reasoning model. To correct the teaching error, it transfers the garment back to its owner to absorb the hard knowledge in the self-study phase. Guided by the transfer parsing, we adjust the position of the transferred garment via STN to prevent distortion. Afterward, we estimate a progressive flow to precisely warp the garment with shape and content correspondences. To ensure warping rationality, we supervise the training of the garment warping model using target shape and warping knowledge from virtual try-on. To better preserve body features in the transfer result, we propose a well-designed training strategy for the arm regrowth task to infer new exposure skin. Experiments demonstrate that our method has state-of-the-art performance compared with other virtual try-on and garment transfer methods in garment transfer, especially for preserving garment texture and body features.